3. A transformer has 500 primary turns and 2200 secondary turns. The input voltage is 120 v and the output current is 15.0 A. What is the output voltage and input current assuming 100 percent efficiency?
Answer #1

Here ,

Np = 500

Ns = 2200

Vp = 120 V

Io = 15 A

for the output voltage

500/2200 = 120/Vs

Vs = 528 V

the output voltage is 528 V

for the input current

120 * Iin = 15 * 528

Iin = 66 A

the input current is 66 A